# Online Multi-Agent Pickup and Delivery with Task Deadlines

### v1

@@ Abstract
Managing delivery deadlines in automated warehouses and factories is crucial for maintaining customer satisfaction and ensuring seamless production. This study introduces the problem of online multi-agent pickup and delivery with task deadlines (MAPD-D), which is an advanced variant of the online MAPD problem incorporating delivery deadlines. MAPD-D presents a dynamic deadline-driven approach that includes task deadlines, with tasks being added at any time (online), thus challenging conventional MAPD frameworks. To tackle MAPD-D, we propose a novel algorithm named deadline-aware token passing (D-TP). The D-TP algorithm is designed to calculate pickup deadlines and assign tasks while balancing execution cost and deadline proximity. Additionally, we introduce the D-TP with task swaps (D-TPTS) method to further reduce task tardiness, enhancing flexibility and efficiency via task-swapping strategies. Numerical experiments were conducted in simulated warehouse environments to showcase the effectiveness of the proposed methods. Both D-TP and D-TPTS demonstrate significant reductions in task tardiness compared to existing methods, thereby contributing to efficient operations in automated warehouses and factories with delivery deadlines.
 [^0] This work has been submitted to the IEEE for possible publication. Copyright may be transferred without notice, after which this version may no longer be accessible.
@@ I Introduction
The multi-agent pickup and delivery (MAPD) problem is an extension of MAPF, wherein paths for multiple items are planned from pickup locations to delivery locations [9, 10, 11]. MAPD can be applied to various environments, including automated warehouses and factories [9, 12, 13]. In such settings, the importance of managing delivery deadlines cannot be overstated. Warehouses must tailor deadlines for individual orders to ensure customer satisfaction, whereas factories require timely deliveries to maintain seamless production. Therefore, satisfying the deadlines appropriately is essential for operational efficiency and economic success.
To consider deadlines in MAPD, we define the online multi-agent pickup and delivery with task deadlines (MAPD-D) as a new problem (Fig. 1). Existing studies [14, 15] have explored MAPD with deadlines in an offline setting, where all task information is provided in advance. However, to the best of our knowledge, no study has addressed deadline-aware MAPD in an online setting where tasks may be added at any time. Hence, this study introduces new problem definitions for MAPD-D.
We propose the deadline-aware token passing (D-TP) algorithm to tackle MAPD-D. This algorithm is designed to calculate pickup deadlines and assign tasks while striking a balance between execution cost and deadline proximity. Additionally, we introduce the D-TP with task swaps (D-TPTS) method to reduce task tardiness. D-TPTS enhances flexibility and efficiency by employing task-swapping strategies among agents and within a single agent.
Figure 1: Simulation examples in a four-neighbor grid environment [9]. The black cells represent blocked areas; dark gray cells indicate task endpoints; and light gray cells denote non-task endpoints. Each agent searches for a path from the pickup location to the delivery location for the assigned task. Unlike MAPD [9], a deadline is set for each task in MAPD-D.
@@ II Related Work
Ma et al. [9] reported that online MAPD instances are solvable if they are well-formed. They introduced locations referred to as endpoints, where agents can remain without blocking other agents. Endpoints include all pickup and delivery locations along with the initial positions and designated parking locations. The pickup and delivery locations are referred to as task endpoints, whereas the other locations serve as non-task endpoints. An MAPD instance is well-formed if and only if (a) the number of tasks is finite, (b) the number of agents is not greater than non-task endpoints, and (c) a path exists between any two endpoints without traversing others.
@@
Several researchers have investigated MAPF and MAPD with deadlines. In this study, we classified related studies from the literature considering four perspectives (Table I). “Multi-task” indicates whether each agent is continuously assigned tasks; “Deadlines” denote whether each task is set with a deadline; “Individual release times” denote whether the release time of each task is the same; and “Online” indicate whether tasks are added at any time. As shown in Table I, Ma et al. [16], Wang and Chen [17], and Huang et al. [18] considered deadlines in the context of MAPF. Wu et al. [14] and Ramanathan et al. [15] introduced deadlines in MAPD. However, the perspective of online tasks in MAPD was not considered in [14, 15]. In this study, the proposed MAPD-D considers online tasks with deadlines, where tasks can be added at any time.
@@ III Problem Definition
In this section, we describe the shared definitions of MAPD [9] and MAPD-D and define the tasks specific to MAPD-D.
An instance of MAPD and MAPD-D consists of $m$ agents in $A={a_{1},\cdots,a_{m}}$, a connected simple undirected graph $G=(V,E)$, and a set of unexecuted tasks $\mathcal{T}={\tau_{1},\cdots,\tau_{k}}$. Here, $V$ represents the vertices, $E$ denotes the edges, and $l_{i}(t)\in V$ indicates the location of agent $a_{i}$ at timestep $t$. A path is a sequence of vertices associated with timesteps, indicating the vertex at which the agent is located at each timestep.
Agents either remain at their current node $l_{i}(t)=l_{i}(t+1)$ or move to an adjacent node via an edge $\left(l_{i}(t),l_{i}(t+1)\right)\in E$ at each timestep. Agents must avoid collisions with each other. MAPD and MAPD-D define two types of collisions [1, 9]: (1) Vertex conflict, where two agents cannot occupy the same location at the same timestep, i.e., for all agents $a_{i},a_{j}\,(i\neq j)$ and all timesteps $t$, $l_{i}(t)\neq l_{j}(t)$ must hold; and (2) Swapping conflict, where two agents cannot move along the same edge in opposite directions at the same timestep, i.e., for all agents $a_{i},a_{j}\,(i\neq j)$ and all timesteps $t$, $l_{i}(t)\neq l_{j}(t+1)$ or $l_{j}(t)\neq l_{i}(t+1)$ must hold.
We further provide an extended definition of tasks in MAPD-D. Each task $\tau_{j}\in\mathcal{T}$ consists of a pickup location $v^{\mathrm{p}}_{j}\in V$, a delivery location $v^{\mathrm{d}}_{j}\in V$, and a delivery deadline $d^{\mathrm{d}}_{j}$. New tasks can be added to the task set $\mathcal{T}$ at each timestep. When an agent assigned to a task reaches the delivery location via the pickup location, the task is completed, and its completion time is denoted as $c_{j}$. Once a task is completed, the agent is assigned to a new task.
The objective of MAPD is to identify paths that execute all tasks in a timely manner, whereas MAPD-D aims to minimize task tardiness. We define the tardiness for the $j$-th task as $\epsilon_{j}=\max(0,c_{j}-d^{\mathrm{d}}_{j})$. For each task $\tau_{j}$, $\epsilon_{j}=0$ indicates the success of the task, whereas $\epsilon_{j}>0$ indicates task failure. The objective function of MAPD-D is selected from the following two options:
@@
    where $U(\cdot)$ is a unit step function  [^1] The unit step function. $U(x)$ =$=\begin{cases}0&\text{if }x\leq 0\\ 1&\text{if }x>0\end{cases}$
@@
In this study, we used the cumulative tardiness (2) as the objective function because it provides a more detailed evaluation of tardiness compared to the objective function (1), which only considers success or failure.
@@ IV Proposed Method
This section outlines the D-TP algorithm employed to address the MAPD-D problem, which is an extension of the existing TP method. Typically, TP assigns tasks to agents solely based on execution cost. To reduce task tardiness, we introduce enhancements in two key areas: the calculation of pickup deadlines and deadline-aware task assignment. Additionally, we present D-TPTS, which improves flexibility and efficiency through task-swapping strategies.
@@ IV-A1 Calculation of Pickup Deadlines
The pickup deadline is calculated based on the delivery deadline when a new task $\tau_{j}$ is added. We prepared a dummy agent, implemented using prioritized path planning [11]. D-TP calculates the path for the dummy agent to depart from the delivery location $v^{\mathrm{d}}_{j}$ at time $d^{\mathrm{d}}_{j}$ and move towards the pickup location $v^{\mathrm{p}}_{j}$ by reversing the timesteps. During the path calculation, the order of the pickup and delivery locations is reversed because the time required for transportation can vary depending on the paths of other agents in the environment. The proposed method searches for a path from the delivery location to the pickup location by reversing the order of time, thus calculating the latest possible path (dummy path) that meets the delivery deadline. The pickup deadline $d^{\mathrm{p}}_{j}$ represents the time obtained by subtracting the length (timesteps) of the dummy path from the delivery deadline $d^{\mathrm{d}}_{j}$.
@@ IV-A2 Deadline-aware Task Assignment
The disparity between the calculated pickup deadline $d^{\mathrm{p}}_{j}$ and the current time $t$ indicates the temporal margin of the deadline. In TP, the system assigns tasks to minimize the execution cost at the moment of assignment. In contrast, in D-TP, the system assigns tasks to agents in a manner that minimizes the weighted sum of the execution cost and the temporal margin relative to the deadline.
@@
where $0\leq\alpha\leq 1$ and $\mathcal{T}^{\prime}$ denotes the set of tasks that can be assigned to agent $a_{i}$. The first term $(d^{\mathrm{p}}_{j}-t)$ denotes the temporal margin for the deadline of task $\tau_{j}$; the second term $h(loc(a_{i}),v^{\mathrm{p}}_{j})$ indicates the h-value from the current location of agent $a_{i}$ to the pickup location of the task, which represents the execution cost; and the parameter $\alpha$ indicates the weight for the urgency of the pickup deadline. When $\alpha=0$, it is equivalent to the existing method [9] that does not take the deadline into account.
@@ IV-A3 Algorithm
Algorithm 1 provides the pseudocode for D-TP; parts that differ from TP [9] are indicated in red. In lines 17–20, we define the UpdatePickupDeadline function. In line 19, a dummy agent is prepared to calculate the dummy path from the delivery location $v^{\mathrm{d}}_{j}$ to the pickup location $v^{\mathrm{p}}_{j}$ using the reversed-path finding function $\mathcal{RP}(dummy,v^{\mathrm{d}}_{j},v^{\mathrm{p}}_{j},token,d^{\mathrm{d}}_{j})$. In line 20, the pickup deadline $d^{\mathrm{p}}_{j}$ is calculated by subtracting the length of the dummy path. Here, $|\mathcal{P}|$ represents the length of the path, that is, the number of steps required to move.
@@
3:    Add all new tasks, if any, to task set $\mathcal{T}$
@@
6:        $\mathcal{T}^{\prime}\leftarrow\{\tau_{j}\in\mathcal{T}\,|\,\text{no other pathin }token\text{ ends in }v^{\mathrm{p}}_{j}\text{ or }v^{\mathrm{d}}_{j}\}$
@@
9:           $\tau_{j^{*}}\leftarrow\underset{\tau_{j}\in\mathcal{T}^{\prime}}{\operatorname{argmin}}\left(\alpha\cdot(d^{\mathrm{p}}_{j}-t)+(1-\alpha)\cdot h(loc(a_{i}),v^{\mathrm{p}}_{j})\right)$
@@
19:        Update $\tau_{j}$’s dummy path in $token$ with $\mathcal{RP}(dummy,v^{\mathrm{d}}_{j},v^{\mathrm{p}}_{j},token,d^{\mathrm{d}}_{j})$
20:        Update $\tau_{j}$’s pickup deadline $d^{\mathrm{p}}_{j}$ with $d^{\mathrm{d}}_{j}-|\mathcal{RP}(dummy,v^{\mathrm{d}}_{j},v^{\mathrm{p}}_{j},token,d^{\mathrm{d}}_{j})|$
@@ IV-B1 Task Swapping Among Agents
In contrast to TPTS, which focuses solely on execution cost considerations, the proposed method modifies this approach to incorporate a weighted sum of the execution cost and temporal margin for deadlines, as expressed in (3).
@@ IV-B2 Task Switching
In other words, the following inequalities should hold simultaneously:
$d^{\mathrm{p}}_{\mathrm{new}}<d^{\mathrm{p}}_{\mathrm{cur}}$ (4)
@@
where $\mathrm{cur}$ denotes the index of the current task of the agent $a_{i}$ and $\mathrm{new}$ represents the index of the new task.
@@ IV-B3 Algorithm
Algorithm 2 provides the pseudocode for D-TPTS. The overall flow mirrors that of TPTS [9], with differences highlighted in red. Lines 5–10 implement task switching. When an agent is en route to the pickup locations, it abandons its current task and accepts a different task if the new task has an earlier pickup deadline and a lower execution cost than the current task. Task swapping is performed in the function GetTask (lines 15–32). As indicated in line 19, the system considers both the temporal margin of the pickup deadline and the weighted sum of the task execution cost.
@@
3:    Add all new tasks, if any, to task set $\mathcal{T}$
@@
16:    $\mathcal{T}^{\prime}\leftarrow\{\tau_{j}\in\mathcal{T}\,|\,\text{no other pathin }token\text{ ends in }v^{\mathrm{p}}_{j}\text{ or }v^{\mathrm{d}}_{j}\}$
@@
19:        $\tau_{j^{*}}\leftarrow\underset{\tau_{j}\in\mathcal{T}^{\prime}}{\operatorname{argmin}}\left(\alpha\cdot(d^{\mathrm{p}}_{j}-t)+(1-\alpha)\cdot h(loc(a_{i}),v^{\mathrm{p}}_{j})\right)$

### v2

@@ Abstract
Managing delivery deadlines in automated warehouses and factories is crucial for maintaining customer satisfaction and ensuring seamless production. This study introduces the problem of online multi-agent pickup and delivery with task deadlines (MAPD-D), an advanced variant of the online MAPD problem incorporating delivery deadlines. In the MAPD problem, agents must manage a continuous stream of delivery tasks online. Tasks are added at any time. Agents must complete their tasks while avoiding collisions with each other. MAPD-D introduces a dynamic, deadline-driven approach that incorporates task deadlines, challenging the conventional MAPD frameworks. To tackle MAPD-D, we propose a novel algorithm named deadline-aware token passing (D-TP). The D-TP algorithm calculates pickup deadlines and assigns tasks while balancing execution cost and deadline proximity. Additionally, we introduce the D-TP with task swaps (D-TPTS) method to further reduce task tardiness, enhancing flexibility and efficiency through task-swapping strategies. Numerical experiments were conducted in simulated warehouse environments to showcase the effectiveness of the proposed methods. Both D-TP and D-TPTS demonstrated significant reductions in task tardiness compared to existing methods. Our methods contribute to efficient operations in automated warehouses and factories with delivery deadlines.
000© 2024 IEEE. Personal use of this material is permitted. Permission from IEEE must be obtained for all other uses, in any current or future media, including reprinting/republishing this material for advertising or promotional purposes, creating new collective works, for resale or redistribution to servers or lists, or reuse of any copyrighted component of this work in other works.
@@ I Introduction
The multi-agent pickup and delivery (MAPD) problem is an extension of MAPF, wherein paths for multiple items are planned from pickup to delivery locations without collisions between agents [9, 10, 11]. MAPD can be applied to various environments, including automated warehouses and factories [9, 12, 13]. In such settings, the importance of managing delivery deadlines cannot be overstated. Warehouses must tailor deadlines for individual orders to ensure customer satisfaction, whereas factories require timely deliveries to maintain seamless production. Therefore, satisfying the deadlines appropriately is essential for operational efficiency and economic success.
Figure 1: Simulation examples in a four-neighbor grid environment [9]. Unlike MAPD [9], a deadline is set for each task in MAPD-D. Each agent searches for a path from the pickup to delivery locations for the assigned task.
Figure 2: Comparison of task assignment methods: Our method assigns tasks by considering the execution cost and proximity of deadlines, thereby effectively reducing tardiness.
To consider deadlines in MAPD, we defined the online multi-agent pickup and delivery with task deadlines (MAPD-D) as a new problem (Fig. 1). Existing studies [14, 15] have explored MAPD with deadlines in an offline setting, where all task information is provided in advance. However, to the best of our knowledge, no study has addressed deadline-aware MAPD in an online setting where tasks may be added at any time. Hence, this study introduces new problem definitions for MAPD-D.
We propose the deadline-aware token passing (D-TP) algorithm to tackle MAPD-D. This algorithm is designed to calculate pickup deadlines and assign tasks while striking a balance between execution cost and deadline proximity (Fig. 2). Additionally, we introduce the D-TP with task swaps (D-TPTS) method to reduce task tardiness. D-TPTS enhances flexibility and efficiency by employing task-swapping strategies among agents and within a single agent.
@@ II Related Work
Several researchers have investigated MAPF and MAPD with deadlines. In this study, we classified related studies from the literature considering four perspectives (Table I). “Multi-task” indicates whether each agent is continuously assigned tasks; “Deadlines” denote whether each task is set with a deadline; “Individual release times” denote whether the release time of each task is the same; and “Online” indicates whether tasks are added at any time. As summarized in Table I, Ma et al. [16], Wang and Chen [17], and Huang et al. [18] considered deadlines in the context of MAPF. Wu et al. [14] and Ramanathan et al. [15] introduced deadlines in MAPD. However, the perspective of online tasks in MAPD was not considered in [14, 15]. In this study, the proposed MAPD-D considers online tasks with deadlines, where tasks can be added at any time.
The dynamic vehicle routing problem with pickups and deliveries (Dynamic VRPPD) also considers deadlines for online requests (tasks) [19, 20]. Passengers’ request time corresponds to deadlines, and one of the objective functions in Dynamic VRPPD is to minimize the sum of the difference between the request and pickup times (waiting time) [21]. These settings are similar to MAPD-D; however, Dynamic VRPPD does not consider collisions between agents.
@@ III Problem Definition
III-A MAPD-D problem
In this sub-section, we describe the shared definitions of MAPD [9] and MAPD-D and define the tasks specific to MAPD-D.
An instance of MAPD and MAPD-D consists of $m$ agents in $A={a_{1},\ldots,a_{m}}$, a connected simple undirected graph $G=(V,E)$, and a set of unexecuted tasks $\mathcal{T}={\tau_{1},\ldots,\tau_{k}}$. Here, $V$ represents the vertices, $E$ denotes the edges, and $l_{i}(t)\in V$ indicates the location of agent $a_{i}$ at timestep $t$. A path is a sequence of vertices associated with timesteps, indicating the vertex at which the agent is located at each timestep.
Agents either remain at their current node $l_{i}(t)=l_{i}(t+1)$ or move to an adjacent node via an edge $\left(l_{i}(t),l_{i}(t+1)\right)\in E$ at each timestep. Agents must avoid collisions with each other. MAPD and MAPD-D define two types of collisions [1, 9]: (1) Vertex conflict, where two agents cannot occupy the same location at the same timestep, that is, for all agents $a_{i},a_{j}\,(i\neq j)$ and all timesteps $t$, $l_{i}(t)\neq l_{j}(t)$ must hold; and (2) Swapping conflict, where two agents cannot move along the same edge in opposite directions at the same timestep, that is, for all agents $a_{i},a_{j}\,(i\neq j)$ and all timesteps $t$, $l_{i}(t)\neq l_{j}(t+1)$ or $l_{j}(t)\neq l_{i}(t+1)$ must hold.
We further provide an extended definition of tasks in MAPD-D. Each task $\tau_{j}\in\mathcal{T}$ consists of a pickup location $v^{\mathrm{p}}_{j}\in V$ and delivery location $v^{\mathrm{d}}_{j}\in V$. In an extended definition for MAPD-D, each task additionally includes a delivery deadline $d^{\mathrm{d}}_{j}$. $d^{\mathrm{d}}_{j}$ is not a tight bound, and agents are still expected to take on new tasks and proceed even if a task’s deadline has passed. When an agent assigned to a task reaches the delivery location via the pickup location, the task is completed, and its completion time is denoted as $c_{j}$. Once a task is completed, the agent is assigned a new task. New tasks can be added to the task set $\mathcal{T}$ at each timestep.
The objective of MAPD is to identify paths that execute all tasks promptly, whereas MAPD-D aims to minimize task tardiness. We define the tardiness for the $j$-th task as $\epsilon_{j}=\max(0,c_{j}-d^{\mathrm{d}}_{j})$. For each task $\tau_{j}$, $\epsilon_{j}=0$ indicates the success of the task, whereas $\epsilon_{j}>0$ indicates task failure. The objective function of MAPD-D is selected from the following two options:
@@ III-A MAPD-D problem
    where $U(\cdot)$ is a unit step function  [^1] The unit step function. $U(x)$ =$=\begin{cases}0&\text{if }x\leq 0\\ 1&\text{if }x>0.\end{cases}$
@@
In this study, we used the cumulative tardiness (2) as the objective function because it provides a more detailed evaluation of tardiness compared with the objective function (1), which only considers success or failure.
III-B Well-formed MAPD-D Instances
Ma et al. [9] reported that online MAPD instances are solvable if they are well-formed. They introduced locations referred to as endpoints, where agents can remain without blocking other agents. Endpoints include all pickup and delivery locations along with the initial positions and designated parking locations. The pickup and delivery locations are referred to as task endpoints, whereas the other locations serve as non-task endpoints. An MAPD instance is well-formed if and only if (a) the number of tasks is finite, (b) the number of agents does not exceed the number of non-task endpoints, and (c) a path exists between any two endpoints without traversing others.
In MAPD-D, only the task deadline is introduced, while other task settings and environments are the same as in MAPD. Therefore, the definition of well-formed MAPD-D instances is the same as that for MAPD.
@@ IV Proposed Method
This section outlines the D-TP algorithm employed to address the MAPD-D problem, which is an extension of the existing TP method. Typically, TP assigns tasks to agents solely based on the execution cost. To reduce task tardiness, we introduce enhancements in two key areas: the calculation of pickup deadlines and deadline-aware task assignments. Additionally, we present D-TPTS, which improves flexibility and efficiency through task-swapping strategies.
@@ IV-A1 Calculation of Pickup Deadlines
The pickup deadline is calculated based on the delivery deadline when a new task $\tau_{j}$ is added. This deadline is not a tight bound but rather a reference value used for cost calculation in subsequent task assignments.
We prepared a dummy agent and implemented it using prioritized path planning [11]. D-TP calculates the path for the dummy agent to depart from the delivery location $v^{\mathrm{d}}_{j}$ at time $d^{\mathrm{d}}_{j}$ and move toward the pickup location $v^{\mathrm{p}}_{j}$ by reversing the timesteps. During the path calculation, the order of the pickup and delivery locations is reversed because the time required for transportation can vary depending on the paths of other agents in the environment. The proposed method searches for a path from the delivery location to the pickup location by reversing the order of time, thus calculating the latest possible path (dummy path) that meets the delivery deadline. The pickup deadline $d^{\mathrm{p}}_{j}$ represents the time obtained by subtracting the length (timesteps) of the dummy path from the delivery deadline $d^{\mathrm{d}}_{j}$.
@@ IV-A2 Deadline-aware Task Assignment
The disparity between the calculated pickup deadline $d^{\mathrm{p}}_{j}$ and current time $t$ indicates the temporal margin of the deadline. In TP, the system assigns tasks to minimize the execution cost at the moment of assignment. Contrastingly, in D-TP, the system assigns tasks to agents in a manner that minimizes the weighted sum of the execution cost and temporal margin relative to the deadline.
@@
where $0\leq\alpha\leq 1$ and $\mathcal{T}^{\prime}$ denotes the set of tasks that can be assigned to agent $a_{i}$. The first term $(d^{\mathrm{p}}_{j}-t)$ denotes the temporal margin for the deadline of task $\tau_{j}$; the second term $h(loc(a_{i}),v^{\mathrm{p}}_{j})$ indicates the h-value from the current location of agent $a_{i}$ to the pickup location of the task, which represents the execution cost; and parameter $\alpha$ indicates the weight for the urgency of the pickup deadline. When $\alpha=0$, it is equivalent to the existing method [9] that does not consider the deadline.
@@ IV-A3 Algorithm
Algorithm 1 provides the pseudocode for D-TP; parts that differ from TP [9] are indicated in blue. In lines 17–20, we define the UpdatePickupDeadline function. In line 19, a dummy agent $a_{dmy}$ is prepared to calculate the dummy path from the delivery location $v^{\mathrm{d}}_{j}$ to the pickup location $v^{\mathrm{p}}_{j}$ using the reversed-path finding function $\mathcal{RP}(dummy,v^{\mathrm{d}}_{j},v^{\mathrm{p}}_{j},token,d^{\mathrm{d}}_{j})$. In line 20, the pickup deadline $d^{\mathrm{p}}_{j}$ is calculated by subtracting the length of the dummy path. Here, $|\mathrm{path}|$ represents the length of the path, that is, the number of steps required to move.
@@
3:    Add all new tasks, if any, to the task set $\mathcal{T}$
@@
6:        $\mathcal{T}^{\prime}\leftarrow\{\tau_{j}\in\mathcal{T}\mid\text{no other path in }token\text{ ends in }v^{\mathrm{p}}_{j}\text{ or }v^{\mathrm{d}}_{j}\}$
@@
9:           $\tau_{j^{*}}$ is calculated by (3)
@@
19:        Update $\tau_{j}$’s dummy path with $\mathcal{RP}(a_{dmy},v^{\mathrm{d}}_{j},v^{\mathrm{p}}_{j},token,d^{\mathrm{d}}_{j})$
20:        Update $\tau_{j}$’s pickup deadline $d^{\mathrm{p}}_{j}$ with $d^{\mathrm{d}}_{j}-|\mathrm{dummy\ path}|$
@@ IV-B1 Task Swapping Among Agents
Contrary to TPTS, which focuses solely on execution cost considerations, the proposed method modifies this approach to incorporate a weighted sum of the execution cost and temporal margin for deadlines, as expressed in (3).
@@ IV-B2 Task Switching
Specifically, the following inequalities should hold simultaneously:
$d^{\mathrm{p}}_{\mathrm{new}}<d^{\mathrm{p}}_{\mathrm{cur}},$ (4)
@@
where $\mathrm{cur}$ denotes the index of the current task of agent $a_{i}$ and $\mathrm{new}$ represents the index of the new task.
@@ IV-B3 Algorithm
Algorithm 2 provides the pseudocode for D-TPTS. The overall procedure mirrors that of TPTS [9], with differences highlighted in blue. Lines 5–10 implement task switching. When an agent is en route to the pickup locations, it abandons its current task and accepts a different task if the new task has an earlier pickup deadline and a lower execution cost than the current task. Task swapping is performed in the function GetTask (lines 15–32). As indicated in line 19, the system considers both the temporal margin of the pickup deadline and the weighted sum of the task execution cost.
@@
3:    Add all new tasks, if any, to the task set $\mathcal{T}$
@@
16:    $\mathcal{T}^{\prime}\leftarrow\{\tau_{j}\in\mathcal{T}\mid\text{no other path in }token\text{ ends in }v^{\mathrm{p}}_{j}\text{ or }v^{\mathrm{d}}_{j}\}$
@@
19:        $\tau_{j^{*}}$ is calculated by (3)
@@
IV-C Completeness
In well-formed MAPD instances, a path exists among the current location of an agent, pickup location, and delivery location. Additionally, the number of tasks is finite. TP eventually assigns a task to each agent, and each agent can execute the task. Therefore, TP solves any well-formed instances.
As discussed in Section III-B, well-formed MAPD-D instances also satisfy the conditions for being well-formed in MAPD. The difference between TP and D-TP is only in the cost calculation, and all other parts are the same. Therefore, D-TP can solve well-formed instances, just like TP in MAPD.
We also introduce task switching in D-TPTS. An agent may abandon its current task when a highly urgent task appears; however, the agent always finishes its current task within a finite time. Therefore, D-TPTS can solve well-formed instances in the same way that TPTS solves them in MAPD.